OpenStudy (anonymous):

How would you solve for theta: R= 3h when R= (v^2*sin2theta)/g and h=(v^2*sin^2theta)/g

OpenStudy (anonymous):

so far I have (v^2*sin2theta)/g = (3v^2*sin^2theta)/2g

OpenStudy (anonymous):

I then put a common denominator of 2g and finally in the end where I find myself stuck is 2sin2theta=3sin^2theta don't know if perhaps I did something wrong and I just don't know what to do now

OpenStudy (anonymous):

Is this a typo, or is h=R? I'm basing this off of your definition in the OP.

OpenStudy (anonymous):

now that I say that though, would it instead be 3R=h? since R is 3X bigger

OpenStudy (anonymous):

In that case, what you want to do is to solve for \(\theta\) in the equation\[\frac{3\sin^2(\theta)v^2}{2g}=\frac{v^2\sin(2\theta)}{g}.\]

OpenStudy (anonymous):

what just happend :o

OpenStudy (anonymous):

but yes that is what I'm trying to figure out

OpenStudy (anonymous):

\[3\sin^2(\theta)=g\sin(2\theta)\]\[\frac{\sin^2(\theta)}{\sin(2\theta)}=\frac{g}{3}\]\[\frac{\sin^2(\theta)}{2\sin(\theta)\cos(\theta)}=\frac{g}{3}\]\[\tan(\theta)=\frac{2}{3}g\]

OpenStudy (anonymous):

I'm curious as to how you got the first part. The 3sin^2=gsin(2theta)

OpenStudy (anonymous):

I just multiplied both sides by \(2g\).

OpenStudy (anonymous):

Then divided both sides by \(v^2\).

OpenStudy (anonymous):

wouldn't the g's "cancel out? As in the value of g would no longer be in the formula after that

OpenStudy (anonymous):

and is it an identity or something? sin2/2sin*cos = tan?

OpenStudy (anonymous):

The \(g\)'s do cancel out: The LHS cancels out completely, whilst the RHS only partially, i.e.,\[\frac{2g}{g}=g.\]I'm also using the identities\[\sin2u=2\sin u\cos u\]and\[\frac{\sin u}{\cos u}=\tan u\]

OpenStudy (anonymous):

ok, I do see how you got the tan; as for the LHS and RHS; sorry, but I'm lost theree. How is it canceled partially?

OpenStudy (anonymous):

\[\frac{3\sin^2(\theta)v^2}{2g}=\frac{v^2\sin(2\theta)}{g}\]\[\frac{2g}{2g}3\sin^2(\theta)v^2=\frac{2g}{g}v^2\sin(2\theta)\]\[3\sin^2(\theta)v^2=gv^2\sin(2\theta)\]\[3\sin^2(\theta)=g\sin(2\theta)\]

OpenStudy (anonymous):

LOL I realized I made a mistake.

OpenStudy (anonymous):

I'm sorry for confusing you.

OpenStudy (anonymous):

You're right, it should be \(2\), not \(g\)!

OpenStudy (anonymous):

lol, it's ok, I was like how is he getting that :o

OpenStudy (anonymous):

still thanks for everything; you explained the problem out which helped a lot!! :D

OpenStudy (anonymous):

Haha, I'm sorry. :3 That was SO silly of me. I will now go and hide my face in shame. LOL

OpenStudy (anonymous):

nah, don't worry, I a few days back I added 170 and 44 and obtained 116. Who should hide there face more? LOL

OpenStudy (anonymous):

LOL I guess we all have our off days then. :)

OpenStudy (anonymous):

Thanks again! :)

OpenStudy (anonymous):

yw :)